# Spatially flat cosmological quantum spacetimes

### v1

@@ 1 Introduction
It has been known for some time that the unitary irreducible representations of $\mathfrak{so}(4,2)$ in the minimal discrete series [11], labeled by $n\in{\mathbb{N}}_{0}$ and sometimes called doubleton representations, can be used to model the fuzzy hyperboloid $H_{n}^{4}$. This is possible because the doubleton representations descend to irreducible representations of $\mathfrak{so}(4,1)$. In a next step, the fuzzy hyperboloid can be used to model covariant quantum spacetimes via certain projections [34, 27, 28]. Among those quantum spacetimes, the ones that are currently best understood are FLRW spacetimes with negative spatial curvature (k=1k=-1) and a Big Bounce geometry [29, 34, 27, 28, 23, 24, 4]. Such spacetimes have a spatial $SO(3,1)$-symmetry.
We recently extended the set of examples by describing a spatially flat ($k=0$) FLRW spacetime with a Big Bang [12]. In the current paper, we generalize this example to describe a much larger class of spatially flat FLRW quantum spacetimes, which have a spatial $E(3)$-symmetry.
All these cosmological quantum spacetimes should be viewed in the context of the IKKT matrix model [19] formulated in $D=10$ dimensions. In particular, the cosmological $k=0$ quantum spacetime discussed in [12] defines an action
@@
Our paper should also be viewed as part of an ongoing program to construct examples of quantum spacetimes and gravity from matrix theory. We are aware of constructions of other cosmological spacetimes [28, 27, 29, 24, 23, 22] in the framework of the IKKT model, as well as fuzzy de Sitter space [8, 13, 14]. For other related work on the origin of quantum spacetime from matrix theory see e.g. [9, 7, 6, 2, 26, 10, 16]. Let us also mention recent work on the polarized IKKT model [15, 21, 25, 20], a deformation of IKKT that may extend the scope of consistent backgrounds.
@@
where $\sim$ indicates approximate equality in the semi-classical regime and where $\Omega$ is the symplectic volume form on $\mathcal{M}$} (which has dimension $2d$). Such quantization maps are not unique, but preferred maps can often be found by requiring their compatibility with some symmetry group, notably for quantized coadjoint orbits such as the fuzzy hyperboloid described in the following. Another way to describe quantization maps is via (quasi-)coherent states $\mathbin{|}x\rangle$ as $\Phi=\int\Omega\phi(x)|x\rangle\langle x|$; see e.g. [30, 32, 34] for an introduction to this framework.
In the case of covariant quantum spacetimes originating from the fuzzy hyperboloid, it turns out that the underlying symplectic space $\mathcal{M}$} is the six-dimensional projective space ${\mathbb{C}P}^{2,1}$ [28, 27], which can be written as a local bundle
@@
Due to the non-trivial bundle structure of the symplectic space, one encounters a problem. Gauge transformations of the matrix model correspond to symplectomorphisms on $\mathcal{M}$} [34] but not necessarily to diffeomorphisms on ${\mathcal{M}}^{1,3}$. However, it may happen that gauge transformations are related to approximate diffeomorphisms on ${\mathcal{M}}^{1,3}$ in regimes where the contribution of the higher spin variables is negligible. The study of such approximate diffeomorphisms on cosmological $k=0$ quantum spacetimes is one of the main points of this paper. Our definitions and results may be extended to other covariant quantum spacetimes.
@@
While the true IKKT model (1.5) has no quadratic "mass" term, it is sometimes necessary to add an $E(3)$-invariant mass term to the action (1.5) by hand to stabilize spacetime. For example, the background discussed in [12] satisfies the IKKT equations with a non-vanishing mass term.
To embed the background $T^{\mu}$, $\mu=0,\dots,3$, describing cosmological quantum spacetimes into the full IKKT model, and to recover gravity [33], one needs to take into account the coupling of $T^{\mu}$ to the matrices describing the extra dimensions as well as their self-coupling. This has been studied to some extent in the case of cosmological k=1k=-1 quantum spacetimes [34, 5, 22]. In the present paper, we will only outline how this may be extended to the case of $k=0$ quantum spacetimes. We expect that the classical stabilization of extra dimensions via $R$ charge discussed in [22] will carry over to the $k=0$ case, while quantum effects coming from the one-loop effective action need to be taken into account so that the matrices defining spacetime lead to a physically reasonable cosmic evolution.
@@ 1.1 Outline of the paper and summary of the main results
Complementing several recent constructions in the case of k=1k=-1 cosmological quantum spacetimes [29, 28, 27, 31, 34, 24, 22, 4, 5, 23], the main aim of this paper is to elaborate on and generalize the example given by the action $S_{0}[\Phi]$ from (1.1), which was discussed in [12].
@@
The background $T^{\mu}$ in (1.1), which we work out in Section 3 and which was partially investigated in the previous paper [12], can be used as a reference background for cosmological $k=0$ quantum spacetimes. We discuss its algebraic structure, which is richer in the present case than it is for cosmological k=1k=-1 quantum spacetimes: there are two distinct but important actions of $E(3)$, while the respective actions of $SO(3,1)$ in the k=1k=-1 case coincide. We introduce appropriate derivations on ${\mathcal{M}}^{1,3}$ that act also on higher-spin valued functions. In complete analogy to the k=1k=-1 case, this derivation is used to relate vector fields on ${\mathcal{M}}^{1,3}$ to tangential vector fields on $H_{n}^{4}$.
@@
In particular, the cosmological $k=0$ spacetimes described with the background $\tilde{T}^{\mu}$ define a FLRW geometry with cosmic scale factor that depends on $\alpha$ and $\beta$. We illustrate the relation between gauge transformation and diffeomorphisms by verifying that certain distinguished, gauge-equivalent backgrounds yield equivalent FLRW line elements in the timelike and covariant gauge. One of them is the background where the dilaton is constant, which is shown to yield a shrinking FLRW spacetime for $k=0$, rather than an expanding spacetime as in the k=1k=-1 case. However, we can obtain a variety of expanding $k=0$ FLRW spacetimes for different $\alpha$ and/or $\beta$. This computation also demonstrates the use and the consistency of local normal coordinates in the higher-spin framework [34].
Finally, we outline in Section 5 how the background $\tilde{T}^{\mu}$ can be incorporated in the full IKKT model. The details of such an incorporation are beyond the scope of the present paper and are left for future research. However, we concisely discuss the general recipe, which is based on a stabilization of dynamical extra-dimensions and corrections to the IKKT equations at one-loop and which should work similarly to the k=1k=-1 case discussed in [5, 2, 22].
In Appendix A, we list a number of pertinent properties of cosmological k=1k=-1 quantum spacetimes that are needed in our construction of cosmological $k=0$ spacetimes. Appendix B contains a derivation of the concrete form of the $SO(3,1)$ generators in various cases.
@@ 2 Fuzzy hyperboloid and cosmological quantum spacetimes
Let us describe how to obtain cosmological quantum spacetimes from matrix models. Within the scope of this paper, such a construction cannot be fully self-contained. We refer to the book [34] for a comprehensive introduction.
@@ 2.1 Fuzzy hyperboloid
In this paper, we will only consider doubleton representations ${\mathcal{H}}_{n}$ of $\mathfrak{so}(4,2)$, which are labeled by $n\in\mathbb{N}_{0}$, and assume that the $M^{ab}$ are given in one of these representations. There is a standard oscillator construction of the doubleton representations, which is described in detail in [34]. In the doubleton representation labeled by $n$, we have the important identity [34]
@@
where we introduced two dependent parameters $r^{2}$, $r>0$, and $R^{2}=r^{2}\frac{n^{2}-4}{4}$, which will introduce a scale on the fuzzy hyperboloid and the cosmological quantum spacetimes. For further details on doubleton representations, we refer to the existing literature [11, 34].
@@
which means that they can be interpreted as non-commutative coordinates describing the four-dimensional fuzzy hyperboloid $H_{n}^{4}$ of radius $R$. Strictly speaking, $R^{2}$ is negative for $n=0,1$ and vanishes for $n=2$. In the minimal case $n=0$, one can nevertheless describe meaningful spacetimes in some late-time regime [23], but to have a global semi-classical structure one should assume $n\gg 0$ [34]. Except for a few instances where we also consider the minimal case $n=0$, we will always assume $n\gg 0$.
@@ 2.1.1 Semi-classical structure
Here, $H^{4}$ is the four-hyperboloid, $S^{2}$ the two-sphere and $\tilde{\times}$ indicates a local bundle structure and that the isometry group $SO(4,1)$ of $H^{4}$ acts non-trivially on the local fiber in a way respecting the bundle map. The internal two-sphere corresponds to the existence of higher-spin degrees of freedom in the model [34, 27, 28].
@@
The algebra ${\mathcal{C}}:={\mathcal{C}}({\mathcal{M}})$ of functions on $\mathcal{M}$} has a decomposition [34]
@@
which satisfies $x_{a}\eth^{a}=0$. We will see later how $\eth^{a}$ serves to define appropriate derivations on the cosmological spacetimes. The trace on $\mathrm{End}({\mathcal{H}}_{n})$ becomes the integral over $\mathcal{M}$} with respect to the symplectic density as in (1.3). In order to have finite kinetic energy and inner products, functions $\phi\in{\mathcal{C}}$ are subject to certain integrability conditions, which we tacitly assume in this work, cf. [34] for details.
@@ 2.2 Cosmological quantum spacetimes
where the volume form $\Omega^{1,3}:=\rho_{M}\textup{d}^{4}x$ on ${\mathcal{M}}^{1,3}$ is the pull-back of the unique $SO(4,1)$-invariant volume form on the 4-hyperboloid $H^{4}\subset{\mathbb{R}}^{4,1}$ spanned by the $x^{a}$ generators. It can be written explicitly as [27, 12]
@@
where $y^{\mu}$ are the coordinates of ${\mathcal{M}}^{1,3}$ as above. Then, defining $\gamma^{\mu\nu}:=\eta_{\alpha\beta}E^{\alpha\mu}E^{\beta\nu}$, one finds [34]
@@ 2.4 Gauge transformations and symplectomorphisms
On the non-commutative level, a gauge transformation on the fuzzy hyperboloid $H_{n}^{4}$ is implemented by a hermitian operator $\Lambda\in\mathrm{End}({\mathcal{H}}_{n})$. Accordingly, a real-valued function $\Lambda\in{\mathcal{C}}({\mathcal{M}})$ defines an infinitesimal symplectomorphism, i.e., an infinitesimal diffeomorphism preserving the symplectic form, on $\mathcal{M}$} [34],
@@
Before describing more general cosmological $k=0$ spacetimes, let us first recall and elaborate on the reference background considered in [12]. Some relevant identities from the k=1k=-1 case are collected in Appendix A.
@@
where $M^{ab}$, $a,b=0,\dots,5$, are the generators of $SO(4,2)$ described in Section 2, were shown to yield the metric $G_{\mu\nu}$ of a spatially flat FLRW spacetime with a Big Bang [34, 12],
@@
The $T^{\mu}$ generators will define a matrix background [^1] These are special cases of the ”pure backgrounds” considered in [34]. in the IKKT model, while the $Y^{\mu}$ generators are considered as (quantized) local coordinate functions on spacetime. The matrices $T^{\mu}$ and $Y^{\mu}$ satisfy the commutation relations [^2] These commutation relations are very similar to those of k=1k=-1 quantum spacetime upon zooming in to a locally flat patch [17].
@@ 3.1 $E(3)$-subalgebra, invariant quantities and Casimir operators
The first Casimir simply follows from the normalization of the $T^{i}$, while the second follows from the identity (A.3). [^3] Note that to derive $C_{2}$ from (A.3), one needs to keep track of signs carefully because the Minkowski metrics relating $X^{0}$ to $X_{0}=-X^{0}$ and $Y^{0}$ to $Y_{0}=-Y^{0}$ do not coincide.

### v2

@@ 1 Introduction
It has been known for some time that the unitary irreducible representations of $\mathfrak{so}(4,2)$ in the minimal discrete series [Fernando:2009fq], labeled by $n\in{\mathbb{N}}_{0}$ and sometimes called doubleton representations, can be used to model the fuzzy hyperboloid $H_{n}^{4}$. This is possible because the doubleton representations descend to irreducible representations of $\mathfrak{so}(4,1)$. In a next step, the fuzzy hyperboloid can be used to model covariant quantum spacetimes via certain projections [Steinacker:2024unq, Sperling:2019xar, Sperling:2018xrm]. Among those quantum spacetimes, the ones that are currently best understood are FLRW spacetimes with negative spatial curvature (k=1k=-1) and a Big Bounce geometry [Steinacker:2017bhb, Steinacker:2024unq, Sperling:2019xar, Sperling:2018xrm, Manta:2025inq, Manta:2024vol, Battista:2022hqn]. Such spacetimes have a spatial $SO(3,1)$-symmetry.
We recently extended the set of examples by describing a spatially flat ($k=0$) FLRW spacetime with a Big Bang [Gass:2025tae]. In the current paper, we generalize this example to describe a much larger class of spatially flat FLRW quantum spacetimes, which have a spatial $E(3)$-symmetry.
All these cosmological quantum spacetimes should be viewed in the context of the IKKT matrix model [Ishibashi:1996xs] formulated in $D=10$ dimensions. In particular, the cosmological $k=0$ quantum spacetime discussed in [Gass:2025tae] defines an action
@@
Our paper should also be viewed as part of an ongoing program to construct examples of quantum spacetimes and gravity from matrix theory. We are aware of constructions of other cosmological spacetimes [Sperling:2018xrm, Sperling:2019xar, Steinacker:2017bhb, Manta:2024vol, Manta:2025inq, MantaStein_deformed] in the framework of the IKKT model, as well as fuzzy de Sitter space [Buric:2017yes, Gazeau:2006hj, Gazeau:2009mi]. For other related work on the origin of quantum spacetime from matrix theory see e.g. [Chaney:2015ktw, Brahma:2022hjv, Brahma:2021tkh, Asano:2024def, Nishimura:2019qal, Chou:2024sgk, Hirasawa:2023lpb]. Let us also mention recent work on the polarized IKKT model [Hartnoll:2024csr, Komatsu:2024ydh, Martina:2025kwc, Komatsu:2024bop, Chou:2025rwy], a deformation of IKKT that may extend the scope of consistent backgrounds.
@@
where $\sim$ indicates approximate equality in the semi-classical regime and where $\Omega$ is the symplectic volume form on $\mathcal{M}$} (which has dimension $2d$). Such quantization maps are not unique, but preferred maps can often be found by requiring their compatibility with some symmetry group, notably for quantized coadjoint orbits such as the fuzzy hyperboloid described in the following. Another way to describe quantization maps is via (quasi-)coherent states $\mathbin{|}x\rangle$ as $\Phi=\int\Omega\phi(x)|x\rangle\langle x|$; see e.g. [Steinacker:2019fcb, Steinacker:2020nva, Steinacker:2024unq] for an introduction to this framework. For the specific case ${\mathcal{M}}\cong{\mathbb{C}}P^{1,2}$ under consideration, these maps are discussed in some detail in Appendix A.
In the case of covariant quantum spacetimes originating from the fuzzy hyperboloid, it turns out that the underlying symplectic space $\mathcal{M}$} is the six-dimensional projective space ${\mathbb{C}P}^{2,1}$ [Sperling:2018xrm, Sperling:2019xar], which can be written as a local bundle
@@
Due to the non-trivial bundle structure of the symplectic space, one encounters a problem. Gauge transformations of the matrix model correspond to symplectomorphisms on $\mathcal{M}$} [Steinacker:2024unq] but not necessarily to diffeomorphisms on ${\mathcal{M}}^{1,3}$. However, it may happen that gauge transformations are related to approximate diffeomorphisms on ${\mathcal{M}}^{1,3}$ in regimes where the contribution of the higher spin variables is negligible. The study of such approximate diffeomorphisms on cosmological $k=0$ quantum spacetimes is one of the main points of this paper. Our definitions and results may be extended to other covariant quantum spacetimes.
@@
While the true IKKT model (1.5) has no quadratic "mass" term, it is sometimes necessary to add an $E(3)$-invariant mass term to the action (1.5) by hand to stabilize spacetime. For example, the background discussed in [Gass:2025tae] satisfies the IKKT equations with a non-vanishing mass term.
To embed the background $T^{\mu}$, $\mu=0,\dots,3$, describing cosmological quantum spacetimes into the full IKKT model, and to recover gravity [Steinacker:2023myp], one needs to take into account the coupling of $T^{\mu}$ to the matrices describing the extra dimensions as well as their self-coupling. This has been studied to some extent in the case of cosmological k=1k=-1 quantum spacetimes [Steinacker:2024unq, Battista:2023glw, MantaStein_deformed]. In the present paper, we will only outline how this may be extended to the case of $k=0$ quantum spacetimes. We expect that the classical stabilization of extra dimensions via $R$ charge discussed in [MantaStein_deformed] will carry over to the $k=0$ case, while quantum effects coming from the one-loop effective action need to be taken into account so that the matrices defining spacetime lead to a physically reasonable cosmic evolution.
@@ 1.1 Outline of the paper and summary of the main results
Complementing several recent constructions in the case of k=1k=-1 cosmological quantum spacetimes [Steinacker:2017bhb, Sperling:2018xrm, Sperling:2019xar, Steinacker:2019awe, Steinacker:2024unq, Manta:2024vol, MantaStein_deformed, Battista:2022hqn, Battista:2023glw, Manta:2025inq], the main aim of this paper is to elaborate on and generalize the example given by the action $S_{0}[\Phi]$ from (1.1), which was discussed in [Gass:2025tae].
@@
The background $T^{\mu}$ in (1.1), which we work out in Section 3 and which was partially investigated in the previous paper [Gass:2025tae], can be used as a reference background for cosmological $k=0$ quantum spacetimes. We discuss its algebraic structure, which is richer in the present case than it is for cosmological k=1k=-1 quantum spacetimes: there are two distinct but important actions of $E(3)$, while the respective actions of $SO(3,1)$ in the k=1k=-1 case coincide. We introduce appropriate derivations on ${\mathcal{M}}^{1,3}$ that act also on higher-spin valued functions. In complete analogy to the k=1k=-1 case, this derivation is used to relate vector fields on ${\mathcal{M}}^{1,3}$ to tangential vector fields on $H_{n}^{4}$.
@@
In particular, the cosmological $k=0$ spacetimes described with the background $\tilde{T}^{\mu}$ define a FLRW geometry with cosmic scale factor that depends on $\alpha$ and $\beta$. We illustrate the relation between gauge transformation and diffeomorphisms by verifying that certain distinguished, gauge-equivalent backgrounds yield equivalent FLRW line elements in the timelike and covariant gauge. One of them is the background where the dilaton is constant, which is shown to yield a shrinking FLRW spacetime for $k=0$, rather than an expanding spacetime as in the k=1k=-1 case. However, we can obtain a variety of expanding $k=0$ FLRW spacetimes for different $\alpha$ and/or $\beta$. This computation also demonstrates the use and the consistency of local normal coordinates in the higher-spin framework [Steinacker:2024unq].
Finally, we outline in Section 5 how the background $\tilde{T}^{\mu}$ can be incorporated in the full IKKT model. The details of such an incorporation are beyond the scope of the present paper and are left for future research. However, we concisely discuss the general recipe, which is based on a stabilization of dynamical extra-dimensions and corrections to the IKKT equations at one-loop and which should work similarly to the k=1k=-1 case discussed in [Battista:2023glw, Asano:2024def, MantaStein_deformed].
In Appendix B, we list a number of pertinent properties of cosmological k=1k=-1 quantum spacetimes that are needed in our construction of cosmological $k=0$ spacetimes. Appendix C contains a derivation of the concrete form of the $SO(3,1)$ generators in various cases.
@@ 2 Fuzzy hyperboloid and cosmological quantum spacetimes
Let us describe how to obtain cosmological quantum spacetimes from matrix models. Within the scope of this paper, such a construction cannot be fully self-contained. We refer to the book [Steinacker:2024unq] for a comprehensive introduction.
@@ 2.1 Fuzzy hyperboloid
In this paper, we will only consider doubleton representations ${\mathcal{H}}_{n}$ of $\mathfrak{so}(4,2)$, which are labeled by $n\in\mathbb{N}_{0}$, and assume that the $M^{ab}$ are given in one of these representations. There is a standard oscillator construction of the doubleton representations, which is described in detail in [Steinacker:2024unq]. In the doubleton representation labeled by $n$, we have the important identity [Steinacker:2024unq]
@@
where we introduced two dependent parameters $r^{2}$, $r>0$, and $R^{2}=r^{2}\frac{n^{2}-4}{4}$, which will introduce a scale on the fuzzy hyperboloid and the cosmological quantum spacetimes. For further details on doubleton representations, we refer to the existing literature [Fernando:2009fq, Steinacker:2024unq].
@@
which means that they can be interpreted as non-commutative coordinates describing the four-dimensional fuzzy hyperboloid $H_{n}^{4}$ of radius $R$. Strictly speaking, $R^{2}$ is negative for $n=0,1$ and vanishes for $n=2$. In the minimal case $n=0$, one can nevertheless describe meaningful spacetimes in some late-time regime [Manta:2025inq], but to have a global semi-classical structure one should assume $n\gg 0$ [Steinacker:2024unq]. Except for a few instances where we also consider the minimal case $n=0$, we will always assume $n\gg 0$.
@@ 2.1.1 Semi-classical structure
Here, $H^{4}$ is the four-hyperboloid, $S^{2}$ the two-sphere and $\tilde{\times}$ indicates a local bundle structure and that the isometry group $SO(4,1)$ of $H^{4}$ acts non-trivially on the local fiber in a way respecting the bundle map. The internal two-sphere corresponds to the existence of higher-spin degrees of freedom in the model [Steinacker:2024unq, Sperling:2019xar, Sperling:2018xrm].
@@
The algebra ${\mathcal{C}}:={\mathcal{C}}({\mathcal{M}})$ of functions on $\mathcal{M}$} has a decomposition [Steinacker:2024unq]
@@
which satisfies $x_{a}\eth^{a}=0$. We will see later how $\eth^{a}$ serves to define appropriate derivations on the cosmological spacetimes. The trace on $\mathrm{End}({\mathcal{H}}_{n})$ becomes the integral over $\mathcal{M}$} with respect to the symplectic density as in (1.3). In order to have finite kinetic energy and inner products, functions $\phi\in{\mathcal{C}}$ are subject to certain integrability conditions, which we tacitly assume in this work, cf. [Steinacker:2024unq] for details.
@@ 2.2 Cosmological quantum spacetimes
where the volume form $\Omega^{1,3}:=\rho_{M}\textup{d}^{4}x$ on ${\mathcal{M}}^{1,3}$ is the pull-back of the unique $SO(4,1)$-invariant volume form on the 4-hyperboloid $H^{4}\subset{\mathbb{R}}^{4,1}$ spanned by the $x^{a}$ generators. It can be written explicitly as [Sperling:2019xar, Gass:2025tae]
@@
where $y^{\mu}$ are the coordinates of ${\mathcal{M}}^{1,3}$ as above. Then, defining $\gamma^{\mu\nu}:=\eta_{\alpha\beta}E^{\alpha\mu}E^{\beta\nu}$, one finds [Steinacker:2024unq]
@@ 2.4 Gauge transformations and symplectomorphisms
On the non-commutative level, a gauge transformation on the fuzzy hyperboloid $H_{n}^{4}$ is implemented by a hermitian operator $\Lambda\in\mathrm{End}({\mathcal{H}}_{n})$. Accordingly, a real-valued function $\Lambda\in{\mathcal{C}}({\mathcal{M}})$ defines an infinitesimal symplectomorphism, i.e., an infinitesimal diffeomorphism preserving the symplectic form, on $\mathcal{M}$} [Steinacker:2024unq],
@@
Before describing more general cosmological $k=0$ spacetimes, let us first recall and elaborate on the reference background considered in [Gass:2025tae]. Some relevant identities from the k=1k=-1 case are collected in Appendix B.
@@
where $M^{ab}$, $a,b=0,\dots,5$, are the generators of $SO(4,2)$ described in Section 2, were shown to yield the metric $G_{\mu\nu}$ of a spatially flat FLRW spacetime with a Big Bang [Steinacker:2024unq, Gass:2025tae],
@@
The $T^{\mu}$ generators will define a matrix background [^1] These are special cases of the ”pure backgrounds” considered in [Steinacker:2024unq]. in the IKKT model, while the $Y^{\mu}$ generators are considered as (quantized) local coordinate functions on spacetime. The matrices $T^{\mu}$ and $Y^{\mu}$ satisfy the commutation relations [^2] These commutation relations are very similar to those of k=1k=-1 quantum spacetime upon zooming in to a locally flat patch [Steinacker:2026qzk].
@@ 3.1 $E(3)$-subalgebra, invariant quantities and Casimir operators
The first Casimir simply follows from the normalization of the $T^{i}$, while the second follows from the identity (B.3). [^3] Note that to derive $C_{2}$ from (B.3), one needs to keep track of signs carefully because the Minkowski metrics relating $X^{0}$ to $X_{0}=-X^{0}$ and $Y^{0}$ to $Y_{0}=-Y^{0}$ do not coincide.